OpenStudy (anonymous):

cos^-1(cos 7pi/6) = ?

OpenStudy (anonymous):

@pooja195 @rational @Miracrown @Zale101 @KingGeorge @sleepyjess @Nnesha

OpenStudy (kinggeorge):

Are you assuming that \(\cos^{-1}\) is the inverse cosine function?

OpenStudy (anonymous):

Yeah!

OpenStudy (kinggeorge):

In that case, the only thing you have to be careful about, is the range of \(\cos^{-1}\). Usually with inverse functions you have that \(f^{-1}(f(x))=x\). But with trig functions it's a bit trickier. So the first step, is to figure out what \(\cos(7\pi/6)\) is. So what is it? How well do you remember your unit circle?

OpenStudy (anonymous):

sec im glitched

OpenStudy (anonymous):

its not letting me post :/

OpenStudy (anonymous):

Not very well, I remember it well for the first quadrant and sort of okayish on the 2nd quadrant. Probably why these problems are harder. But I do know [0, pi] is the principle value branch for cos and that cos^-1(cos 7pi/6) would give 7pi/6?

rishavraj (rishavraj):

hey \[\cos 7\pi/6 = \cos(\pi + \pi/6)\]

rishavraj (rishavraj):

proceed further

OpenStudy (anonymous):

Why take cos 7pi/6 as cos(pi + pi/6)?

rishavraj (rishavraj):

remember?? \[\cos^{-1} (\cos \theta) = \cos \theta~for~all~~ \theta \epsilon[0 , \pi ]\]

OpenStudy (anonymous):

Yeah I do! :)

OpenStudy (anonymous):

But cos^-1 and cos cancel each other and we get theta only?

OpenStudy (kinggeorge):

What rishavraj said. The problem is that \(7\pi/6\) isn't in that interval. So we can't just cancel like you want to.

rishavraj (rishavraj):

here its 7pi/6 \[7\pi/6 > \pi\]

OpenStudy (anonymous):

Oh so we get 7pi/6 into a form which belongs to the principle value branch of cos^-1

rishavraj (rishavraj):

yeah now just apply cos (pi + x) =????

OpenStudy (anonymous):

cos( pi + x) = cos x well actually 2pi + x

rishavraj (rishavraj):

\[\cos (\pi + x) = -\cos x\]

OpenStudy (anonymous):

sec writing that down, my bad

OpenStudy (anonymous):

Oh my god. Lovely pdf, I missed so many! I'll get that printed and copy it down, learn them, thank you O_O

OpenStudy (kinggeorge):

While I do like that approach, I think it may not be the best approach as it requires the memorization of a bunch of extra identities that aren't really necessary...

rishavraj (rishavraj):

yeah i know only few r important .....its kinda rubbish......lol

OpenStudy (anonymous):

Just when I thought I'm finally done with this chapter and its collection of identities, more come flying at me xD

OpenStudy (kinggeorge):

If you want to try (what I think is) an easier approach, we need to know what \(\cos(7\pi/6)\) is.

rishavraj (rishavraj):

@KingGeorge what u think cos7pi/6 = i m blank.....

OpenStudy (anonymous):

The back of the book says the answer should be 5pi/6

OpenStudy (kinggeorge):

And that's the answer since \(\cos(7\pi/6)=\cos(5\pi/6)\). So in theory if you can figure out what \(\cos(7\pi/6)\) is all you need to do is match it up with an \(x\in[0,\pi]\) such that \(\cos(x)=\cos(7\pi/6)\).

rishavraj (rishavraj):

@kennybm its correct........

OpenStudy (anonymous):

ugh this is frustrating me so much. 2pi - 7pi/6?